OpenStudy (anonymous):

How many liters of methane gas (CH4) need to be combusted to produce 12.4 liters of water vapor, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) yields CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

I did this : 12.4 g H2O x (1 mol CH4 divided by 2 mol H2O = 6.20 L CH4

OpenStudy (anonymous):

I just wanna know If i did it right

OpenStudy (anonymous):

@jordanloveangel @jeffyblood @kathlyn98 @KatieMTB @Luigi0210 @Librarian @LiteNing1337 @LittleRedRiding @NathalyN @omgitsjc @Ondinana @OpenStudyRocks5* @pooja195 @raynagarcia @sammixboo @Squirrels @shaleiah @SimonN1995

OpenStudy (anonymous):

Pleaz?

OpenStudy (anonymous):

@TheSmartOne @taylor12344 @Thatkid_marc @thelaw2 @vera_ewing @xo.minnie.xox @xstreetprowl @YOUNGSUPERMAN @zzr0ck3r

OpenStudy (anonymous):

im not good at chemi

OpenStudy (anonymous):

I got the same answer, but at the beginning H20 is in liters, not grams, so you have to use stoichiometry.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I can show the work I did if you want

OpenStudy (anonymous):

yes pleaz

OpenStudy (anonymous):

stiociometry is really hard for me, Sorry

OpenStudy (anonymous):

*stoichiometry

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh I got the second problem done its: You know moles = volume/molar volume (Vm) n(H2O) = 12.4/(Vm) n(CH4)=n(H2O) * 0.5 = 6.2/Vm V(CH4) = 6.4/Vm *Vm = 6.2L

OpenStudy (allieeslabae):

Do you still need help. Is this a certain curriculum...?

OpenStudy (anonymous):

why did u X it by 0.5?

OpenStudy (anonymous):

and it's Chemistry

OpenStudy (anonymous):

@thelaw2

OpenStudy (ondinana):

Sorry aye, I'm not very great in chemistry. All language arts and history is my thing. :P

OpenStudy (anonymous):

@cecormier I think that @thelaw2 is using a different formula that they have memorized, from physics perhaps? But if you are supposed to be using stoichiometry, I'm pretty sure that my work is correct.

OpenStudy (anonymous):

Thank you :)